# Testing New Physics in Oscillations at a Neutrino Factory

### v1

@@ Abstract
A neutrino factory is a potential successor to the upcoming generation of neutrino oscillation experiments and a possible precursor to next-generation muon colliders. Such a machine would provide a well characterized beam of $\nu_{\mu}$, $\bar{\nu}_{\mu}$, $\nu_{e}$, and $\bar{\nu}_{e}$ neutrinos with comparable statistics. Here we show the sensitivity of a neutrino factory to new oscillation physics scenarios such as vector neutrino non-standard interactions and CPT violation. We study two different potential setups for a neutrino factory with different assumptions on charge identification in the far detector. We find that a neutrino factory can improve over most of the current constraints on these scenarios and over forecasted constraints by DUNE, even when doubling DUNE’s statistics. Additionally, we find that a neutrino factory can break degeneracies between the standard oscillation parameters and neutrino non-standard interaction parameters present at DUNE.
@@ 1 Introduction
However, there exist degeneracies between many of the new physics parameters and the standard oscillation parameters, which may make it challenging to search for new physics even with the upcoming planned experiments. There are several experimental proposals beyond the upcoming JUNO, Hyper-Kamiokande, and DUNE programs that aim to break these degeneracies and improve the sensitivities to both BSM parameters and the standard oscillation parameters. First, a multi-baseline setup such as the T2HKK experiment Hyper-Kamiokande:2016srs , long-baseline experiments targetting the second maximum such as ESSnuSB ESSnuSB:2021azq , and atmospheric neutrino experiments Razzaque:2014vba ; Kelly:2019itm can resolve the degeneracies by measuring neutrino oscillations at different baselines. Secondly, similar to the DUNE experiment DUNE:2020ypp the P2O experiment Akindinov:2019flp can search for differences between BSM and standard oscillations in a broad energy spectrum. Thirdly, an additional low-energy neutrino source from muon decay at rest Ge:2016dlx can reduce the new physics impact and provide a clean measurement of standard oscillation parameters, which in turn gives a better sensitivity to new physics parameters when combined with a high-energy neutrino beam. Finally, performing a joint analysis of all experiments is desirable to resolve degeneracies and benefit from complementarities to provide stringent constraints on BSM scenarios (as done in Coloma:2023ixt in the case of new interactions).
@@
In this study, we consider another possibility for a potential future neutrino oscillation experiment, namely a neutrino factory (NF) which uses LAr far detectors to measure high-energy $\mathcal{O}$(GeV) $\overset{\scriptscriptstyle(-)}{\nu}_{\mu}$ and $\overset{\scriptscriptstyle(-)}{\nu}_{e}$ beams coming from muon decays Geer:1997iz ; Blondel:2000gj ; Blondel:2006su ; FernandezMartinez:2010zza . Recently, standard three-flavor oscillations at a modern neutrino factory setup have been studied Denton:2024glz as well as non-oscillation new physics in Bogacz:2022xsj . In this paper we study new physics in oscillations at a neutrino factory. We focus on vector neutrino non-standard interaction (NSI) and CPT violation where long-baseline experiments can provide the strongest bounds. Other BSM scenarios, like scalar NSI Ge:2018uhz ; Babu:2019iml are much better constrained by solar neutrino experiments due to the significant amplification of this effect inside the sun Denton:2024upc and the sterile neutrino scenario is typically best constrained by short-baseline experiments Machado:2019oxb ; Dentler:2018sju ; Boser:2019rta and atmospheric neutrinos Nunokawa:2003ep ; IceCubeCollaboration:2024nle . A neutrino factory mainly benefits from three perspectives to constrain NSI and CPT violation compared to existing and planned oscillation experiments, like DUNE or Hyper-Kamiokande, which use neutrinos produced in a fixed target setup: i) a larger statistics of the oscillation channel from the electron flavor to the electron/muon flavor makes it sensitive to more new physics parameters, ii) the higher neutrino energy can amplify the new physics effects such as NSI to make it easier to be observed, and iii) the simultaneous presence of equally large amounts of neutrinos and antineutrinos, albeit with a different energy spectrum, in the beam allows for studies of new physics in the neutrino and anti-neutrino sectors separately.
@@
The paper is outlined as follows. In Sec. 2, we discuss the experimental setup of a neutrino factory, in sec. 3, 4 we present our results for the the scenarios of new neutrino interactions, and CPT violation before we conclude in sec. 5.
@@ 2 Experimental Setup
At a neutrino factory an intense neutrino beam with $\mathcal{O}$(GeV) energy can be produced via the decays of positively (negatively) charged muons inside a muon storage ring. Due to the 3-body muon decay process, the flavor composition and energy spectrum of the neutrinos are well understood. Moreover, different from traditional accelerator neutrino experiments, there exists a large fraction of $\nu_{e}$ ($\bar{\nu}_{e}$) from $\mu^{+}$ ($\mu^{-}$) decay which allows studies of $\nu_{e}\rightarrow\nu_{e},\nu_{\mu}$ ($\bar{\nu}_{e}\rightarrow\bar{\nu}_{e},\bar{\nu}_{\mu}$) oscillation channels with large statistics. We consider two possible setups for a neutrino factory, motivated by existing accelerator complexes in the US. [^4] See Kitano:2024kdv for a study of a neutrino factory based at J-PARC. Either the neutrinos are produced at i) Fermilab (FNAL) or at ii) Brookhaven National Laboratory (BNL) Denton:2024glz . We assume $10^{21}$ muon decays per year with equal running time for muons and anti-muons. For the FNAL-SURF (BNL-SURF) baseline we use as maximal neutrino energy $E_{\mu}=5(8)$ GeV. The choice of energy and running time is motivated by previous results from Denton:2024glz as these values optimize the precision of the CP phase $\delta$ in the absence of new physics at a neutrino factory; for different goals such as new physics one may prefer different muon energies. We assume that the far detector is the DUNE far detector with a total of 40 kt LAr fiducial mass to detect neutrinos produced at FNAL or BNL, which have traveled a baseline of $L=1284.9\,$km or $L=2542.3\,$km, respectively. For both configurations we use a matter density of 2.8482.8482.848\,2.848g/cm3 and we assume a 2% uncertainty on matter density which is consistent with DUNE TDR DUNE:2021cuw . Realistically, the matter density for the BNL configuration is somewhat higher due to its longer baseline however, similarly to DUNE, a neutrino factory is expected to be sensitive to the matter density at the 30% level Kelly:2018kmb , thus a slightly different value due to the longer baseline will not significantly affect our results.
@@
The beam at a neutrino factory has a large intrinsic backgrounds, for example in the $\nu_{e}$ disappearance channel coming from $\bar{\nu}_{\mu}\to\bar{\nu}_{e}$ appearance. It is possible to differentiate them at some level due to the different initial spectrum combined with the (presumably) different oscillation probability. The excellent energy reconstruction abilities of LArTPC can partially differentiate these components to the far detector flux Huber:2008yx and additionally a detector with charge identification (CID) capabilities allows for improved background rejection. To study the role of CID we consider three scenarios of the detector performance: i) no charge identification, ii) with 100% charge identification of electrons ($e$CID), and iii) with 100% charge identification of muons ($\mu$CID). We also include the oscillated tau neutrinos with the charged-current detection process at the far detector. Due to the large uncertainties on the tau neutrino signal efficiency at LAr detectors, however, we treat the tau neutrino events as background only, using the cross section from Yaeggy:2023wfe . We consider only normalization uncertainties at the far detector for a neutrino factory where we use 2.5% for both the muon and electron neutrino normalization uncertainties. For a neutrino factory the statistics of the $\nu_{e},~{}\nu_{\mu}$ in the FD is comparable, therefore we treat the normalization uncertainty for a neutrino factory slightly different than for LBNF neutrinos DUNE:2021cuw , where the muon and electron neutrino signal normalization uncertainties are 5% and 2%, respectively, while a common 5% uncertainty is used for muon and electron neutrino background. Moreover, we take a 10% uncertainty on the neutral-current background, 20% uncertainty on the tau background. We use the GLoBES framework Huber:2004ka ; Huber:2007ji for our studies.
@@
To compare the potential of a NF to planned future experiments we also take DUNE into account [^5] We note that there has been proposed a high energy tune for DUNE designed to enhance its sensitivity for tau neutrinos, NSIs, and other scenarios Masud:2017bcf ; Masud:2018pig ; DeGouvea:2019kea ; Rout:2020cxi ; Ghoshal:2019pab ; Siyeon:2024pte , but would do so at the cost of shifting away from the CP optimized beam. The NF beam parameters we have chosen here were selected to maximize the sensitivity to CP violation Denton:2024glz . Thus for a direct comparison, we focus on DUNE in the nominal configuration., where the experimental details can be found in DUNE:2020ypp ; DUNE:2021cuw . We consider the following experimental setups: i) 10 years of DUNE (DUNE-10), ii) 20 years of (DUNE-20) [^6] We assume $1.1\times 10^{21}$ POT per year of DUNE DUNE:2021cuw ., or iii) 10 years of DUNE plus 10 years of a neutrino factory at either FNAL (NF-FNAL) and BNL (NF-BNL) baseline. Each of the two NF scenarios has three CID scenarios as mentioned above. Therefore we have eight different experimental setups in total.
@@ 3 Neutrino non-standard Interactions
We first introduce the framework for (vector) NSI and then show the sensitivity of DUNE and a NF to the NSI parameters and the standard oscillation parameters in the presence of NSI. We focus on $\delta_{CP}^{\text{true}}=-90^{\circ}$, the results for $\delta_{CP}^{\text{true}}=0^{\circ}$ can be found in the appendix.
@@ 3.1 Framework
Neutrinos could have interactions with other fermions due to the presence of new mediators. A general framework to incorporate this effect is neutrino non-standard interactions which, for example, originate from dimension-6 operators. First proposed by Wolfenstein Wolfenstein:1977ue , NSIs are often considered to have a vector mediator. In fact, such new neutrino interactions are natural features in many neutrino mass models with new mediators Ohlsson:2012kf ; Miranda:2015dra ; Babu:2019mfe . The strength of the NSI depends on the specific model and large NSI is predicted in some models, see e.g. Forero:2016ghr ; Denton:2018dqq ; Dey:2018yht ; Babu:2017olk ; Farzan:2016wym ; Farzan:2015hkd ; Farzan:2015doa ; Babu:2019mfe . In general, vector NSI can be of charged-current (CC) or neutral-current (NC) type Farzan:2017xzy ; Proceedings:2019qno . Both affect neutrinos during their propagation through matter while only CC NSI affects neutrinos at the point of production and detection. [^8] Note that NC NSI does affect the detection of neutrinos in NC channels, see e.g. Coloma:2017egw ; Gehrlein:2024vwz . Long-baseline experiments like DUNE and a NF are strongly affected by matter effects, and thus are expected to have exceptional sensitivity to the NC NSI. To this end, we take the NC NSI into consideration below in the context of propagation. [^9] One could also consider NC NSI in the detection of the NC channel at a NF.
In the low-energy regime, neutrino NC NSI with matter fields can be formulated in terms of the effective four-fermion Lagrangian term Denton:2022nol ,
@@ 3.2 Results for non-standard vector interactions
Figure 1: Constraints on the diagonal (left) NSI parameters and off-diagonal NSI parameters (taken to be real) (right) at 90% confidence level assume 10 years of DUNE (red), 20 years of DUNE, and 10 years of DUNE combined with 10 years of a NF located at FNAL or BNL (hatched). The purple, yellow, and green color correspond to no CID at the NF far detector, 100% electron CID, and 100% $\mu$ CID. We compare the results to the global fit results from Coloma:2023ixt . There are three allowed regions of $\epsilon_{ee}-\epsilon_{\mu\mu}$ from the global fit, where two of them are shown by error bars while the remaining one $[-3.1,-2.8]$ is shown with numbers.
@@
We start with the simplest case, where all NSI parameters are assumed to be real and only one NSI parameter is considered at one time. While this will miss some of the more interesting physics cases presented below, it is valuable in highlighting the relative role of each experiment. In Fig. 1, we show the expected constraints on the diagonal and the real part of the off-diagonal NSI parameters at the 90% confidence level (C.L.) in the left and right panel, respectively. As discussed in Sec. 2, there are eight experimental scenarios under consideration, the allowed regions are shown by patches of different colors. For comparison, we also include the current global fit data Coloma:2023ixt with an error bar. Note however that for the global fit results the NSI parameters not shown have been marginalized over, in contrast to our approach.
@@
We find that for all five NSI parameters, DUNE combined with a neutrino factory can reach better sensitivities than DUNE alone, even with doubled running time. Especially for a neutrino factory at BNL with $e$/$\mu$CID, the sensitivity can be improved by a factor of $\sim 2$ compared to DUNE with 20 years running time. We also generally find that it is possible to improve beyond global bounds even in cases where non-accelerator experiments are expected to dominate. The first of which is $\epsilon_{ee}-\epsilon_{\mu\mu}$ which is strongly constrained by the combination of solar and reactor data, and yet DUNE and a NF may be better than existing constraints. The second of which are the two parameters tightly constrained in atmospheric experiments: $\epsilon_{\tau\tau}-\epsilon_{\mu\mu}$ and $\epsilon_{\mu\tau}$. We find that DUNE and a NF can actually outperform the existing constraints on $\epsilon_{\mu\tau}$. Finally, we note that in terms of the NSI parameters expected to be best probed by long-baseline accelerator experiments, $\epsilon_{e\mu}$ and $\epsilon_{e\tau}$, we find that DUNE with a NF can outperform, in some cases easily, existing constraints.
@@
There are two separate regions in the $\epsilon_{ee}-\epsilon_{\mu\mu}$ parameter space due to the LMA-Dark solution Miranda:2004nb ; Escrihuela:2009up ; Bakhti:2014pva ; Coloma:2016gei ; Denton:2021vtf . This is a degeneracy in the oscillation Hamiltonian in matter in the presence of NSI. The Hamiltonian in the case of standard matter effect is physically equivalent to the Hamiltonian $H$ in the presence of NSI by taking $\Delta m^{2}_{21}\to-\Delta m^{2}_{21}$, $\Delta m^{2}_{31}\to-\Delta m^{2}_{31}$, $\delta_{CP}\to-\delta_{CP},$ $\epsilon_{ee}-\epsilon_{\mu\mu}=-2$, $\epsilon_{\tau\tau}-\epsilon_{\mu\mu}=0$, and $\epsilon_{\alpha\beta}\in\mathbb{R}$.
@@
Figure 2: Expected constraints on different combinations of two non-zero NSI parameters at 90% confidence level. All NSI parameters are assumed to be real, that is, the complex phase is $\in\{0,\pi\}.$ We compare the expected constraints from DUNE only (with 10 and 20 years of running time) and DUNE combined with a NF for different assumptions on CID. Note that the $\epsilon_{e\mu}$-$\epsilon_{e\tau}$ panel has different axes.
@@
To study the degeneracies between different NSI parameters, we extend our study to the two-parameter space which we anticipate should cover the majority of the degeneracy space. We first consider the case where the off-diagonal NSI parameters are assumed to be real and the experimental constraints at 90% C.L. are shown in Fig. 2. There exists a significant degeneracy between $\Re\epsilon_{e\mu}$ and $\Re\epsilon_{e\tau}$, see e.g. Liao:2016hsa ; Chatterjee:2021wac , which is because the only relevant combinations of NSI parameters in the $\nu_{e}/\bar{\nu}_{e}$ appearance channel are $\epsilon_{e\mu}+\epsilon_{e\tau}$ and $\epsilon_{e\mu}-\epsilon_{e\tau}$, assuming real NSI parameters Denton:2022pxt . Especially when $\delta_{CP}=-90^{\circ}$, only the combination of $\epsilon_{e\mu}-\epsilon_{e\tau}$ survives Denton:2022pxt , which leads to the positive correlation between $\epsilon_{e\mu}$ and $\epsilon_{e\tau}$ as shown in the figure. In all cases except the $\epsilon_{\tau\tau}-\epsilon_{\mu\mu}$-$\epsilon_{ee}-\epsilon_{\mu\mu}$ panel, DUNE combined with a neutrino factory at BNL with $\mu$CID configuration gives the most stringent constraints on the two-parameter space due to the longer baseline of the BNL setup. In the $\epsilon_{\tau\tau}-\epsilon_{\mu\mu}$-$\epsilon_{ee}-\epsilon_{\mu\mu}$ panel, the $e$CID and $\mu$CID cases are both very similar and both provide the strongest constraint.

### v2

@@ Abstract
A neutrino factory is a potential successor to the upcoming generation of neutrino oscillation experiments and a possible precursor to next-generation muon colliders. Such a machine would provide a well-characterized beam of $\nu_{\mu}$, $\bar{\nu}_{\mu}$, $\nu_{e}$, and $\bar{\nu}_{e}$ neutrinos with comparable statistics. Here we show the sensitivity of a neutrino factory to new oscillation physics scenarios such as vector neutrino non-standard interactions and CPT violation. We study two different potential setups for a neutrino factory with different assumptions on charge identification in the far detector. We find that 10 years of a neutrino factory combined with 10 years of DUNE can improve over most of the current constraints on these scenarios and even over forecasted constraints by 20 years of DUNE. Additionally, we find that a neutrino factory can break degeneracies between the standard oscillation parameters and neutrino non-standard interaction parameters present at DUNE.
@@ 1 Introduction
However, there exist degeneracies between many of the new physics parameters and the standard oscillation parameters, which may make it challenging to search for new physics even with the upcoming planned experiments. There are several experimental proposals beyond the upcoming JUNO, Hyper-Kamiokande, and DUNE programs that aim to break these degeneracies and improve the sensitivities to both BSM parameters and the standard oscillation parameters. First, a multi-baseline setup such as the proposed long-baseline experiments T2HKK Hyper-Kamiokande:2016srs and ESSnuSB ESSnuSB:2021azq targeting the second maximum could provide a different handle on the oscillation physics, as could the proposed P2O experiment Akindinov:2019flp . Second, an additional low-energy neutrino source from muon decay at rest Ge:2016dlx can reduce the new physics impact and provide a clean measurement of standard oscillation parameters, which in turn gives a better sensitivity to new physics parameters when combined with a high-energy neutrino beam. Finally, performing a joint analysis of all experiments is desirable to resolve degeneracies and benefit from complementarities to provide stringent constraints on BSM scenarios (as done in Coloma:2023ixt in the case of new interactions).
@@
In this study, we consider another possibility for a potential future neutrino oscillation experiment, namely a neutrino factory (NF) which uses LAr far detectors to measure high-energy $\mathcal{O}$(GeV) $\overset{\scriptscriptstyle(-)}{\nu}_{\mu}$ and $\overset{\scriptscriptstyle(-)}{\nu}_{e}$ beams coming from muon decays Geer:1997iz ; DeRujula:1998umv ; Blondel:2000gj ; Albright:2000xi ; Blondel:2006su ; FernandezMartinez:2010zza . Recently, standard three-flavor oscillations at a modern neutrino factory setup were studied in Denton:2024glz as well as non-oscillation new physics in Bogacz:2022xsj . In this paper we study new physics in oscillations at a neutrino factory. We focus on vector neutrino non-standard interaction (NSI) and CPT violation where long-baseline experiments can provide the strongest bounds for many of the oscillation parameters. Other BSM scenarios, like scalar NSI Ge:2018uhz ; Babu:2019iml are much better constrained by solar neutrino experiments due to the significant amplification of this effect inside the sun Denton:2024upc and the sterile neutrino scenario is typically best constrained by short-baseline experiments Machado:2019oxb ; Dentler:2018sju ; Boser:2019rta and atmospheric neutrinos Nunokawa:2003ep ; IceCubeCollaboration:2024nle . A neutrino factory mainly benefits from three perspectives to constrain NSI and CPT violation compared to existing and planned oscillation experiments, like DUNE or Hyper-Kamiokande, which use neutrinos produced in a fixed target setup: i) a larger statistics of the oscillation channels from the electron flavor to the electron and muon flavors makes it sensitive to more new physics parameters, ii) the higher neutrino energy can amplify the new physics effects such as NSI allowing for an easier observation of their effects, and iii) the simultaneous presence of equally large amounts of neutrinos and anti-neutrinos, albeit with a different energy spectrum, in the beam allows for studies of new physics in the neutrino and anti-neutrino sectors separately. Vector NSIs have been studied in the context of a NF or similar experiments in the past, see e.g. Huber:2001zw ; Huber:2002bi ; Apollonio:2002en ; Blennow:2005qj ; ISSPhysicsWorkingGroup:2007lul ; Kopp:2007mi ; Kopp:2008ds ; Meloni:2009cg ; Antusch:2009pm ; Bakhti:2016prn and CPT violation at a neutrino factory has been studied in Bilenky:2001ka ; Antusch:2008zj ; this study makes use of an updated perspective on the oscillation parameters, the global constraints on new physics, and an up-to-date picture of the upcoming experimental landscape.
@@
The paper is outlined as follows. In section 2, we discuss the experimental setup of a neutrino factory. In sections 3 and 4 we present our results for the the scenarios of new neutrino interactions and CPT violation. We conclude in sec. 5.
@@ 2 Experimental Setup
At a neutrino factory an intense neutrino beam with $\mathcal{O}$(GeV) energy can be produced via the decays of positively (negatively) charged muons inside a muon storage ring. Due to the 3-body muon decay process, the flavor composition and energy spectrum of the neutrinos are well understood. Moreover, different from traditional accelerator neutrino experiments, there exists a large fraction of $\nu_{e}$ ($\bar{\nu}_{e}$) from $\mu^{+}$ ($\mu^{-}$) decay which allows for studies of $\nu_{e}\rightarrow\nu_{e},\nu_{\mu}$ ($\bar{\nu}_{e}\rightarrow\bar{\nu}_{e},\bar{\nu}_{\mu}$) oscillation channels with large statistics. We consider two possible setups for a neutrino factory, motivated by existing accelerator complexes in the US. [^4] See Kitano:2024kdv for a study of a neutrino factory based at J-PARC. Either the neutrinos are produced at i) Fermilab (FNAL) or at ii) Brookhaven National Laboratory (BNL) Denton:2024glz . We assume $10^{21}$ muon decays per year ISSDetectorWorkingGroup:2007uvu ; Geer:2010zz with equal running time for muons and anti-muons. For the FNAL-SURF (BNL-SURF) baseline we use as maximal neutrino energy $E_{\mu}=5(8)$ GeV. The choice of energy and running time is motivated by previous results from Denton:2024glz as these values optimize the precision of the CP phase $\delta$ in the absence of new physics at a neutrino factory; for different goals such as new physics one may prefer different muon energies. We assume that the far detector is the DUNE far detector with a total of 40 kt LAr fiducial mass DUNE:2020lwj to detect neutrinos produced at FNAL or BNL, which have traveled a baseline of $L=1284.9\,$km or $L=2542.3\,$km, respectively. For both configurations we use a matter density of 2.8482.848\,2.848g/cm3 and we assume a 2% uncertainty on matter density which is consistent with the DUNE TDR DUNE:2021cuw . Realistically, the matter density for the BNL configuration is somewhat higher due to its longer baseline however, similarly to DUNE, a neutrino factory is expected to be sensitive to the matter density at the 30% level Kelly:2018kmb , thus a slightly different value due to the longer baseline will not significantly affect our results.
The beam at a neutrino factory has a large intrinsic background, for example in the $\nu_{e}$ disappearance channel coming from $\bar{\nu}_{\mu}\to\bar{\nu}_{e}$ appearance. It is possible to differentiate the two channels at some level due to the different initial spectrum combined with the (presumably) different oscillation probability. The excellent energy reconstruction abilities of LArTPC detectors can partially separate these components at the far detector Huber:2008yx and additionally a detector with charge identification (CID) capabilities allows for improved background rejection. To study the role of CID we consider three scenarios of the detector performance: i) no charge identification, ii) with 100% charge identification of electrons ($e$CID), and iii) with 100% charge identification of muons ($\mu$CID). Existing studies have proposed a variety of possible CID scenarios for the DUNE far detectors. For example, for muon neutrino CID, DUNE could differentiate a muon final state from an anti-muon final state with $\epsilon_{CID}\approx 72\%$ Ternes:2019sak from muon capture on argon. In addition, DUNE may get some CID information from inelasticity as well. In Huber:2008yx it has been pointed out that CID can also be achieved statistically at some level, provided that the energy resolution is sufficiently good. Electron CID at GeV energies has been studied in the context of a magnetized LAr detector Rubbia:1977zz ; Rubbia:2001pk ; Rubbia:2004tz ; Rubbia:2009md . Other detector configurations may achieve significantly improved CID capabilities; given that the final detector details for DUNE are still under consideration, such a discussion is extremely timely.
@@
We also include the oscillated tau neutrinos with the charged-current detection process at the far detector. Due to the large uncertainties on the tau neutrino signal efficiency at LAr detectors, however, we treat the tau neutrino events as background only, using the cross section from Yaeggy:2023wfe . We consider only normalization uncertainties at the far detector for a neutrino factory where we use 2.5% for both the muon and electron neutrino normalization uncertainties because for a neutrino factory the statistics of $\nu_{e},~\nu_{\mu}$ in the FD are comparable. This differs from the approach used for LBNF neutrinos DUNE:2021cuw , where the muon and electron neutrino signal normalization uncertainties are taken to be 5% and 2%, respectively, while a common 5% uncertainty is used for muon and electron neutrino background as the muon neutrino measurement is used to inform the electron neutrino flux. We also take a 10% uncertainty on the neutral-current background and a 20% uncertainty on the tau background. We use the GLoBES framework Huber:2004ka ; Huber:2007ji for our studies.
@@
To compare the potential of a NF to planned future experiments, we also take DUNE into account [^5] We note that there has been a high energy tune for DUNE proposed designed to enhance its sensitivity for tau neutrinos, NSIs, and other new physics scenarios Masud:2017bcf ; Masud:2018pig ; DeGouvea:2019kea ; Rout:2020cxi ; Ghoshal:2019pab ; Siyeon:2024pte . However this advantage comes at the cost of shifting away from the CP optimized beam. The NF beam parameters we have chosen here were also selected to maximize the sensitivity to CP violation Denton:2024glz . Thus for a direct comparison, we focus on DUNE in the nominal configuration., where the experimental details can be found in DUNE:2020ypp ; DUNE:2021cuw . We consider the following experimental setups: i) 20 years of (DUNE-20) [^6] We assume $1.1\times 10^{21}$ POT per year of DUNE DUNE:2021cuw ., or ii) 10 years of DUNE plus 10 years of a neutrino factory at either FNAL (NF-FNAL) and BNL (NF-BNL) baseline. Each of the two NF scenarios has three CID scenarios as mentioned above. Therefore we have seven different experimental setups in total. To keep the plots for the results clean and to limit the number of their curves, we show the results for the BML configuration in the appendix.
@@ 3 Neutrino Non-Standard Interactions
We first introduce the framework for vector NSI and then show the sensitivity of DUNE and a NF to the NSI parameters and the standard oscillation parameters in the presence of NSI. We focus on $\delta_{CP}^{\text{true}}=-90^{\circ}$, the results for $\delta_{CP}^{\text{true}}=0^{\circ}$ can be found in the appendix.
@@ 3.1 Framework
Neutrinos could have interactions with other fermions due to the presence of new mediators. A general framework to incorporate this effect is neutrino non-standard interactions which, for example, originate from dimension-6 operators. First proposed by Wolfenstein Wolfenstein:1977ue , NSIs are often considered to have a vector mediator. In fact, such new neutrino interactions are natural features in many neutrino mass models with new mediators Ohlsson:2012kf ; Miranda:2015dra ; Babu:2019mfe . The strength of the NSI depends on the specific model and large NSI is predicted in some models, see e.g. Forero:2016ghr ; Denton:2018dqq ; Dey:2018yht ; Babu:2017olk ; Farzan:2016wym ; Farzan:2015hkd ; Farzan:2015doa ; Babu:2019mfe . In general, vector NSI can be of charged-current (CC) or neutral-current (NC) type Farzan:2017xzy ; Proceedings:2019qno . Both affect neutrinos during their propagation through matter while only CC NSI affects neutrinos at the point of production and detection. Note also that NC NSI does affect the detection of neutrinos in NC channels, see e.g. Coloma:2017egw ; Gehrlein:2024vwz . Long-baseline experiments like DUNE and a NF are strongly affected by matter effects, and thus are expected to have exceptional sensitivity to the NC NSI. To this end, we take the NC NSI into consideration below in the context of propagation. [^8] One should also consider NC NSI in the detection of the NC channel at a NF. We leave such an analysis at a NF for future work.
In the low-energy regime, neutrino NC NSI with matter fields can be formulated in terms of the effective four-fermion Lagrangian term Wolfenstein:1977ue , (see Antusch:2008tz ; Biggio:2009nt ; Ohlsson:2012kf ; Miranda:2015dra ; Farzan:2017xzy ; Proceedings:2019qno for recent reviews),
@@ 3.2 Results for non-standard vector interactions
Figure 1: Constraints on the diagonal (left) NSI parameters and off-diagonal NSI parameters (taken to be real) (right) at 90% confidence level assume 20 years of DUNE (blue) and 10 years of DUNE combined with 10 years of a NF located at FNAL. The purple, yellow, and green color correspond to different levels of CID at the NF far detector: no CID, 100% electron CID, and 100% $\mu$ CID, respectively. We compare the results to the global fit results from Coloma:2023ixt . There are three allowed regions of $\epsilon_{ee}-\epsilon_{\mu\mu}$ from the global fit, where two of them are shown by error bars while the remaining one [3.1,2.8][-3.1,-2.8][ - 3.1 , - 2.8 ] is shown with numbers. See Fig. 11 in the appendix for the sensitivities of the BNL configuration.
@@
We start with the simplest case, where all NSI parameters are assumed to be real and only one NSI parameter is considered at a time; we will generalize this assumption later. While this will miss some of the more interesting physics cases presented below, it is valuable in highlighting the relative role of each experiment. In Fig. 1, we show the expected constraints on the diagonal and the real part of the off-diagonal NSI parameters at the 90% confidence level (C.L.) in the left and right panels, respectively. We focus on four of the experimental scenarios discussed in Sec. 2. The allowed regions are shown by patches of different colors. For comparison, we also include the current global fit data Coloma:2023ixt with an error bar. Note however that for the global fit results the NSI parameters not shown have been marginalized over.
@@
We find that for all five NSI parameters, DUNE combined with a neutrino factory can reach better sensitivities than DUNE alone, even with doubled running time. As we show in the appendix, especially for a neutrino factory at BNL with $e$/$\mu$CID, the sensitivity can be improved by a factor of 2$\sim$ 2∼ 2 compared to DUNE with 20 years running time. We also generally find that it is possible to improve beyond global bounds even in cases where non-accelerator experiments are expected to dominate. The first of which is $\epsilon_{ee}-\epsilon_{\mu\mu}$ which is strongly constrained by the combination of solar and reactor data, and yet DUNE and a NF may be better than existing constraints. The second of which are the two parameters tightly constrained in atmospheric experiments: $\epsilon_{\tau\tau}-\epsilon_{\mu\mu}$ and $\epsilon_{\mu\tau}$. We find that DUNE and a NF can actually outperform the existing constraints on $\epsilon_{\mu\tau}$ from a recent global fit Coloma:2023ixt as well as newer similar constraints directly from the experiments, see KM3NeT:2024pte . Finally, we note that in terms of the NSI parameters expected to be best probed by long-baseline accelerator experiments, $\epsilon_{e\mu}$ and $\epsilon_{e\tau}$, we find that DUNE with a NF can outperform, in some cases easily, existing constraints.
@@
There are two separate regions in the $\epsilon_{ee}-\epsilon_{\mu\mu}$ parameter space due to the LMA-Dark solution Miranda:2004nb ; Escrihuela:2009up ; Bakhti:2014pva ; Coloma:2016gei ; Denton:2021vtf ; Denton:2022nol . This is an exact degeneracy in the oscillation Hamiltonian in matter in the presence of NSI which requires scattering experiments to break it, see e.g. Coloma:2017egw ; Coloma:2017ncl ; Chaves:2021pey ; Denton:2022nol . The Hamiltonian in the case of the standard matter effect is physically equivalent to the Hamiltonian $H$ in the presence of NSI by taking $\Delta m^{2}_{21}\to-\Delta m^{2}_{21}$, $\Delta m^{2}_{31}\to-\Delta m^{2}_{31}$, $\delta_{CP}\to-\delta_{CP},$ $\epsilon_{ee}-\epsilon_{\mu\mu}=-2$, $\epsilon_{\tau\tau}-\epsilon_{\mu\mu}=0$, and $\epsilon_{\alpha\beta}\in\mathbb{R}$.
@@
Figure 2: Expected constraints on different combinations of two non-zero NSI parameters at 90% confidence level. All NSI parameters are assumed to be real, that is, the complex phase is $\in\{0,\pi\}$. We compare the expected constraints from 20 years of DUNE only and 10 years of DUNE combined with 10 years of NF-FNAL for different assumptions on CID. Note that the $\epsilon_{e\mu}$-$\epsilon_{e\tau}$ panel has different axes. See Fig. 12 in the appendix for the sensitivities for the BNL configuration.
@@
To study other partial degeneracies among different NSI parameters, we extend our study to the two-parameter space which we anticipate should cover the majority of the degeneracy space. We first consider the case where the off-diagonal NSI parameters are assumed to be real and the experimental constraints at 90% C.L. are shown in Fig. 2. There exists a significant degeneracy between $\Re\epsilon_{e\mu}$ and $\Re\epsilon_{e\tau}$, see e.g. Liao:2016hsa ; Chatterjee:2021wac , which is because the dominant combinations of NSI parameters in the $\nu_{e}/\bar{\nu}_{e}$ appearance channel are $\epsilon_{e\mu}+\epsilon_{e\tau}$ and $\epsilon_{e\mu}-\epsilon_{e\tau}$, assuming real NSI parameters Denton:2022pxt . Especially when $\delta_{CP}=-90^{\circ}$, only the combination of $\epsilon_{e\mu}-\epsilon_{e\tau}$ survives Denton:2022pxt , which leads to the positive correlation between $\epsilon_{e\mu}$ and $\epsilon_{e\tau}$ as shown in the figure. As we show in the appendix, in all cases except the $\epsilon_{\tau\tau}-\epsilon_{\mu\mu}$-$\epsilon_{ee}-\epsilon_{\mu\mu}$ panel, DUNE combined with a neutrino factory at BNL with $\mu$CID configuration gives the most stringent constraints on the two-parameter space due to the longer baseline of the BNL setup. In the $\epsilon_{\tau\tau}-\epsilon_{\mu\mu}$-$\epsilon_{ee}-\epsilon_{\mu\mu}$ panel, the $e$CID and $\mu$CID cases are both very similar and both provide strong constraints.